3. If the price elasticity of demand for tractors priced between $4,000 and $6,000 is
-0.75 (using the mid-point method), what will be the percentage change in quantity
demanded when the price for tractors falls from $6,000 to $4,000?
1
Expert's answer
2020-05-25T09:24:39-0400

"Es=\frac{\%∆Q}{\%∆P}"

%∆P="\frac{P2-P1}{\frac{P2+P1}{2}}" =$2000/$5000= 0.4

Es= -0.75

%∆Q=-0.75*0.4= -0.3